OpenStudy (anonymous):

Rewrite sin x^4 so that it involves only the first power of cosine.

OpenStudy (agent0smith):

\[\large \sin ^4 x = (\sin^2 x)(\sin^2 x)\]use the fact that \[\Large \sin^2 x = \frac{ 1-\cos 2x }{ 2}\]

OpenStudy (anonymous):

@agent0smith what is the name of that rule?